OpenStudy (anonymous):

functions f(x)=x^2-x+2,find a) f(0) , b).f(-1), c) f(x+h), and d). (f(x+h)-f(x))/h

OpenStudy (phi):

f(0) means replace x with 0 in the equation, and then do the arithmetic. Try it.

OpenStudy (anonymous):

phi is that the only thing am to do there?

OpenStudy (phi):

that is (a)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and the same thing in b?

OpenStudy (phi):

what did you get for (a) and (b)?

OpenStudy (anonymous):

i got 2 for both questions

OpenStudy (phi):

Now (c), replace x with (x+h) in the equation, and expand things: (x+h)^2 - (x-h) +2 = x^2 +2xh+h^2 -x +h +2

OpenStudy (phi):

Now (d). f(x+h) is as above. f(x) is the original equation. Subtract them: x^2 +2xh+h^2 -x +h +2 - (x^2-x+2)= ?? and then divide your answer by h.

OpenStudy (anonymous):

am hooked,can u help me expand f(x+h)

OpenStudy (anonymous):

?

OpenStudy (phi):

first, you know f(x+h)= (x+h)^2 - (x-h) +2 (all x's replace with (x+h) ) I already expanded it: (x+h)^2 = x^2 +2xh+h^2 and -(x-h) is -x +h so it becomes x^2 +2xh+h^2-x +h+2 Unless you are asking how to expand (x+h)^2 ?

OpenStudy (anonymous):

ok don't worry i was thinking we had to simply further from there

OpenStudy (anonymous):

so that's my answer for c?

OpenStudy (phi):

yes

OpenStudy (anonymous):

cool thanks alot

OpenStudy (anonymous):

hey what is the value of my h in dis case,when i want to divide

OpenStudy (phi):

For (d) you have: f(x+h) - f(x) = x^2 +2xh+h^2 -x +h +2 - (x^2-x+2) this simplifies to 2xh +h +h^2 now, as long as h is not zero, we can divide to get \( \frac{2xh+h+h^{2}}{h}= 2x+1 +h \) In calculus, we let h approach zero (very,very close), so the expression approaches 2x+1

OpenStudy (phi):

\[\frac{2xh+h+h^2}{h}= 2x+1+h\]